OpenStudy (anonymous):

Jeannie is 5 years older than her younger sister April. Together, Jeannie and April’s ages add to 41. How old is April?

OpenStudy (anonymous):

41

OpenStudy (anonymous):

Thats not an answer choice A. 18 B. 23 C. 36 D. 46

OpenStudy (tyteen4a03):

Set up an equation. You know the age of April is x, then Jeanie's age would be x + 5, since she is older.

OpenStudy (anonymous):

So X+5=41 I got C?

OpenStudy (alekos):

J + A = 41 J = A +5

OpenStudy (alekos):

now substitute the second equation into the first and solve for Aprils age

OpenStudy (anonymous):

So this is how i did it. A+5=41 A+5-5=41-5 A=36

OpenStudy (tyteen4a03):

@victoria_lion No, J = A + 5, so the equation is (A + 5) + A = 41.

OpenStudy (alekos):

so then 2A = 36

OpenStudy (alekos):

do you follow?

OpenStudy (anonymous):

Well tell me how c is not the answer @tyteen4a03 ?

OpenStudy (tyteen4a03):

@victoria_lion Because your equation is not correct. See my and @alekos 's explanation above.

OpenStudy (alekos):

without using any algebra you can just take each option given, add 5, and then add the two numbers obtained.

OpenStudy (alekos):

so if we take option A we have 18 + (18+5) = 41, the total of their ages

OpenStudy (anonymous):

Omg okay thats what i was missing i forgot to add their ages *face palm*

OpenStudy (alekos):

and using algebra we also have 2A = 36 so A = 18

OpenStudy (alekos):

i'm glad it clicked

OpenStudy (anonymous):

Thank you so much. :)

OpenStudy (alekos):

no problem

OpenStudy (anonymous):

Do you mind helping me on one last problem?

OpenStudy (alekos):

yeah sure

OpenStudy (anonymous):

thanks The graph below illustrates 2 intersecting lines. If a new line is drawn so that it passes through the red line at y = 0 and the blue line at x = –1, what is the slope of the new line? http://media.education2020.com/evresources/2003-09-02-00-00_files/i0320000.jpg

OpenStudy (alekos):

the two points in question are (-4,0) and (-1,-2) so the new line goes through these points happy with that?

OpenStudy (anonymous):

Haha the way you say i feel stupid XD but yes i am. Ur awesome thanks

OpenStudy (anonymous):

-2/3 right?

OpenStudy (alekos):

no that's ok. i'm just glad that you're now able to see how to solve this problem. you're absolutely right, well done

OpenStudy (anonymous):

Okay thanks :) bye

OpenStudy (alekos):

see you later

OpenStudy (anonymous):

its 36